OpenStudy (anonymous):

Solve: (3k + 11)^2 = 7 here's what I have so fare, please check and/or correct: (3k + 11)(3k + 11) = 7 9k^2 + 33k + 33k + 121 = 7 9k^2 + 66k = 7 - 121 9k^2 + 66k = -1124

OpenStudy (mertsj):

Just leave it in its original form and take the square root of both sides.

OpenStudy (anonymous):

You subtracted wrong. 7 - 121 = -114... Then get it back to the other side and factor. All numbers are divisible by 3 so start there

OpenStudy (anonymous):

Answers are -4.548584 and -2.78475

OpenStudy (anonymous):

okay so answer the first respons i would : the squareroot of 3k + 11 = the square root of 7?

OpenStudy (mertsj):

Yes. Now you're on the right track. Just remember that there are 2 square roots of 7...the positive one and the negative one.

OpenStudy (mertsj):

I should look like this: \[3k+11=\pm \sqrt{7}\]

OpenStudy (mertsj):

Now subtract 11 from both sides:

OpenStudy (anonymous):

divide by 3 on both sides?

OpenStudy (mertsj):

\[3k=-11\pm \sqrt{7}\]

OpenStudy (mertsj):

Now divide by 3. yes.

OpenStudy (anonymous):

so the answer is k = -11 +- the square root of 7/ 3

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

\[k=\frac{-11\pm \sqrt{7}}{3}\]

OpenStudy (anonymous):

okay cool thank u

OpenStudy (mertsj):

yw